OpenStudy (anonymous):

3(5x-25)=15(x-25) is this equation an identity or does it have no solution

OpenStudy (anonymous):

Do the algrebra. Simplify as much as possible and solve for x. There are three possible outcomes. 1) If you get something like x = ? that means that you've found the one possible solution for x. 2) If you get something true like 1=1 0=0 10=10 Then the equation is an identity, meaning it will be true no matter what x is. If you get something false like 1=2 0=1 -10 = 10 Then the equation has no solution. It won't be true for any value of x.

OpenStudy (anonymous):

Try the algebra on your own. If you get stuck, tell me where you're stuck and I'll try ho help explain.

OpenStudy (anonymous):

i got no solution

OpenStudy (anonymous):

Can you tell me why you say that?

OpenStudy (anonymous):

i broke it down and ended with 0=-300

OpenStudy (anonymous):

Exactly right. So this equation is false no matter what x is.

OpenStudy (anonymous):

thanks what about this work problem a path for a new city will connect teh park entrance to main street the path shoud be perpendicular to main street what is an equation to represent the path x is 1 and slope is up 2 over 1

OpenStudy (anonymous):

Does this problem have a picture?

OpenStudy (anonymous):

yes park entrance is on y axis at 5 and main street goes thru x at 1 and slope is up 2 over right 1

OpenStudy (anonymous):

y=-1\2x+5

OpenStudy (anonymous):

 Like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh, hey. You already wrote the equation. Good job. Yes, so the easiest way to write the equation of the new street would be slope intercept form. We need to know two things, the slope and the intercept. The slope comes from the face that it's perpendicular to Main Street. Since perpendicular lines have opposite reciprocal slopes, we get Slope = \(\Large \frac{-1}{2}\) The y-intercept is given. It obviously crosses the y-axis at 5.

OpenStudy (anonymous):

Therefore, like you said, the equation is y = (-1/2)x + 5

OpenStudy (anonymous):

what is the difference between discrete and continous for graph

OpenStudy (anonymous):

Discrete only has specific points, but they won't always be connected. Continuous is a line, meaning it has infinitely many points and they're all connected.  Haha okay, so, first let me say welcome to Open Study. You're a good student. Patient and willing to try problems on you own first. That's great, keep it up. Now let me tell you a couple of things about the site. It is considered most polite, if you have multiple questions, to post them as separate questions instead of commenting on your original question. It's also considered polite to give medals to the best answer. I don't actually care at all about medals, so it doesn't matter in this case, but in the future it will probably result in you getting more help if you do these things.

OpenStudy (anonymous):

Oh, and if you're posting the new problem, and you want to get someone's attention who was just helping you, you can type @ and then start to type their username. It'll let you tag them and they'll get a notification. Like this @bilbr1